OpenStudy (anonymous):

x^2-5x-24=0 quadratic equation...explain while solving?

OpenStudy (anonymous):

FActor it first. Factors of 24 that add up to 5 are 8 and 3. So, you have (x-8)(x+3) = 0. So, value of x can be either 8 or -3.

OpenStudy (anonymous):

x^2 + 5x - 24 = 0 (factor) (x + 8)(x - 3) = 0 (zero-product principle) x + 8 = 0 or x - 3 = 0 x = -8 or x = 3 <===ANSWER

OpenStudy (anonymous):

@GT explain more!